23. Find the molarity and percent mass of the Mg(OH)2 solution using the information below. (12 pts.) For 2 extra credit points, calculate the pH and pOH. Volume of Mg(OH)2 solution Mass of evaporating dish and Mg(OH)2 solution 50.391 g Mass of evaporating dish Mass of evaporating dish and dried Mg(OH)2 10.00 mL 40.380 g 40.868 g
Answer #1

Answer

Molarity of Mg(OH)2 = 0.8367M

Percent mass of Mg(OH)2 = 4.875M%

pOH = 0.22

pH = 3.78

Explanation

Mass of evoparating dish = 40.380g

Mass of evoparating dish + Dried mass of Mg(OH)2 = 40.868g

Difference in mass = 40.868g + 40.380g = 0.488g

mass of Mg(OH)2 = 0.488g

molar mass of Mg(OH)2 = 58.320g/mol

No of mole of Mg(OH)2= 0.488g/58.320g/mol= 0.008367mol

Volume of solution = 10ml

Molarity of Mg(OH)2 = (0.008367mol/10ml)× 1000ml = 0.8367M

[OH-] = 2×0.8367M = 1.6734M

pOH = 0.22

pH = 14 - pOH

pH = 14 - 0.22 = 13.78

Mass of evoparating dish + Mg(OH)2 solution = 50.391g

Mass of evoparating dish = 40.380g

Mass of Mg(OH)2 solution = 10 011

Mass of Mg(OH)2 = 0.488g

Percent mass of Mg(OH)2 = (0.488g/10.011g)×100 = 4.875%